ELECTRONIC TRANSITIONS OF PALLADIUM MONOBORIDE AND PLATINUM MONOBORIDE

\underline{Y. W. NG}, H. F. PANG, Y. S. WONG, YUE QIAN, and A. S-C. CHEUNG , Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.

Electronic transition spectrum of palladium monoboride (PdB) and platinum (PtB) monoboride have been studied using the technique of laser-ablation/reaction free jet expansion and laser induced fluorescence spectroscopy. The metal monoborides were produced by reacting laser ablated metal atoms and diborane ((B_{2}H_{6}) seeded in argon. Five and six vibrational bands were observed respectively for the PdB and PtB molecules. Preliminary analysis of the rotationally resolved structure showed that both molecules have X^{2}Σ^{+} ground state. Least-squares fit of the measured line positions yielded molecular constants for the electronic states involved. Molecular and electronic structures of PdB and PtB are discussed using a molecular orbital energy level diagram. Financial support from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. HKU 701008P) is gratefully acknowledged.